# Novel Node Category Detection Under Subpopulation Shift

### v1

@@ Abstract
In real-world graph data, distribution shifts can manifest in various ways, such as the emergence of new categories and changes in the relative proportions of existing categories. It is often important to detect nodes of novel categories under such distribution shifts for safety or insight discovery purposes. We introduce a new approach, Recall-Constrained Optimization with Selective Link Prediction (RECO-SLIP), to detect nodes belonging to novel categories in attributed graphs under subpopulation shifts. By integrating a recall-constrained learning framework with a sample-efficient link prediction mechanism, RECO-SLIP addresses the dual challenges of resilience against subpopulation shifts and the effective exploitation of graph structure. Our extensive empirical evaluation across multiple graph datasets demonstrates the superior performance of RECO-SLIP over existing methods.
@@ 1 Introduction
Distribution shifts may occur in real-world graphs either through natural temporal evolution [5] or the manual integration of new data sources [30]. Such shifts manifest in various ways, including the emergence of new categories or alterations in the relative proportions of existing ones. The task of novel category detection [3, 27] involves identifying data samples that do not fit into pre-existing categories even in the face of these shifts. As a motivation, consider the product co-purchasing network of a consumer-to-consumer (c2c) e-commerce platform. The co-purchasing network may evolve due to the e-commerce platform being available in new regions. The evolution can lead to the introduction of new product categories and changes in the relative popularity of existing ones. It is crucial to identify products in potential novel categories due to either safety reasons or the need to develop new insights for platform improvement. Additionally, the detection has to be able to be done under the subpopulation shift among existing (non-novel) categories, ideally without knowing the category labels of individual products due to the chaotic nature of a c2c platform. Similar phenomena and demands also exist in academic citation graphs where papers demonstrating new research topics are to be detected. These applications underscore the importance of studying novel node category detection amidst subpopulation shifts.
@@ 2.1 PU Learning and Novel Category Detection
Positive-unlabeled (PU) learning [1] is the problem of learning from a dataset with only positively labeled and unlabeled data. Novel category detection [3, 27] can be naturally reframed as a PU learning problem. The concept of novel and non-novel in our problem can be mapped to the concept of negative and positive in PU learning, respectively. Then, the role of the source nodes would correspond to the positively labeled samples in PU learning since all source nodes are known to be from the non-novel (positive) categories. The target nodes would correspond to the unlabeled samples because they could be from a novel (negative) or non-novel (positive) category and the learner does not know which ones are novel and which ones are not. This connection has been drawn by several prior work [3, 10, 27]. Mainstream PU learning methods [7, 8, 17, 38] deal with the absence of negative labels through risk estimator design, treating unlabeled samples as negatives and labeled samples as weighted combinations of positives and negatives. These risk estimators assume the class priors are given. In practice, the class priors have to be estimated from data by mixture proportion estimation (MPE) techniques [11, 35]. These PU learning approaches are based on the Selected Completely at Random (SCAR) assumption [8]. In our context, SCAR would assume that every node from a non-novel category has an equal probability of being in the source domain.
@@ 4.1 Recall-Constrained Optimization
Figure 2: An illustration of RECO-SLIP. The upper-right module is the recall-constrained optimization component where the classifier adjusts its scores to minimize the FPR on the source while reserving enough target nodes as novel. The bottom module is the selective link prediction component where the link prediction loss is imposed on the target subgraph excluding the nodes with high scores. A solid orange arrow pair between two nodes denotes their representation similarity is maximized whereas a bidirectional hollow arrow denotes the similarity is minimized.
@@ 4.2 Selective Link Prediction
Using the domain labels as a proxy of novel/non-novel labels essentially views all target nodes as equally novel. The novel and non-novel node representations within the target domain will not be sufficiently separated since the domain labels confuse them as the same. However, the edge connection pattern can reveal additional information due to the homophily property. For instance, the target novel nodes should have few edge connections to the target non-novel nodes since they belong to different categories. Therefore, we use link prediction with the graph autoencoder (GAE) [15] objective as an auxiliary task to preserve the category subgroup structure in the node representation space.
@@
$\mathcal{L}_{lp}(g)=-\frac{1}{|\mathcal{E}_{+}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{+}}\log\sigma(\mathbf{g}_{v_{i}}\cdot\mathbf{g}_{v_{j}})-\frac{1}{|\mathcal{E}_{-}^{*}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{-}^{*}}\log(1-\sigma(\mathbf{g}_{v_{i}}\cdot\mathbf{g}_{v_{j}}))$
@@ 5.1.1 Data
We evaluate RECO-SLIP and baseline methods on five widely used public benchmark datasets: Cora [24], CiteSeer [24], Computers [25], Photo [25], and arxiv [13]. Cora, CiteSeer, and arxiv are academic citation graphs while Computers and Photo are product co-purchasing networks. We use the original labels provided by the datasets as the category labels. For Cora, CiteSeer, Computers, and Photo, we view the last class as the novel category and simulate subpopulation shifts by splitting nodes in each category into source and target based on Table 1. We name the preprocessed versions of the four datasets with “-S” at the end representing shift. arxiv is a subset of the ogbn-arxiv graph [13] containing $6$ robotics-related categories from $1990$ to $2012$. arxiv enables a natural source/target split and a novel category emergence since each node is associated with a timestamp and nodes belonging to the category “cs.SY: systems and control” did not exist until $2007$. Therefore, we define nodes with timestamps before $2007$ as the source and the others as the target. In terms of train/test data split, we consider the commonly used transductive node classification setup. The test nodes of arxiv are the nodes with timestamps in $2012$ while the test nodes of the other four datasets are randomly selected from the target domain. We further split a part of the training nodes into the validation set for model selection. We show the dataset-specific statistics in Table 2.
@@ 5.1.2 Baselines
  • SAR-EM [2]: SAR-EM is a propensity-weighting approach for PU-learning when SCAR does not hold. It jointly learns the PU classifier and the propensity score of each data point via the EM algorithm. We select this method as the representative method of propensity-weighting approaches due to its relatively flexible assumptions and publicly available code implementation.
@@ 5.1.3 Hyper-Parameters
All methods except LP-PUL use a two-layer graph convolution network (GCN) [16] coupled with a two-layer MLP as the classifier architecture. We use the Adam optimizer [14] with a learning rate of $0.001$ for both the primal and dual optimizers of RECO-SLIP and all baseline methods excluding LP-PUL. For hyper-parameters specific to RECO-SLIP, we set $\xi=0.001$, $T=1000$, and $\tilde{\bm{\alpha}}=[0.05,0.1,0.15,0.2,0.25]$ for all datasets. By default, we set $\tilde{\beta}$ to $0.01$. However, since the empirical FPR estimate under all $\tilde{\alpha}$ could not go below $0.01$ on Photo-S and CiteSeer-S due to dataset characteristics, we set $\tilde{\beta}$ to $0.05$ for these two datasets.

### v2

@@ Abstract
In real-world graph data, distribution shifts can manifest in various ways, such as the emergence of new categories and changes in the relative proportions of existing categories. It is often important to detect nodes of novel categories under such distribution shifts for safety or insight discovery purposes. We introduce a new approach, Recall-Constrained Optimization with Selective Link Prediction (RECO-SLIP), to detect nodes belonging to novel categories in attributed graphs under subpopulation shifts. By integrating a recall-constrained learning framework with a sample-efficient link prediction mechanism, RECO-SLIP addresses the dual challenges of resilience against subpopulation shifts and the effective exploitation of graph structure. Our extensive empirical evaluation across multiple graph datasets demonstrates the superior performance of RECO-SLIP over existing methods. The experimental code is available at: https://github.com/hsinghuan/novel-node-category-detection.
@@ 1 Introduction
Distribution shifts may occur in real-world graphs either through natural temporal evolution [5] or the manual integration of new data sources [30]. Such shifts manifest in various ways, including the emergence of new categories or alterations in the relative proportions of existing ones. The task of novel category detection [3, 27] involves identifying data samples that do not fit into pre-existing categories even in the face of these shifts. As a motivation, consider the product co-purchasing network of a consumer-to-consumer (c2c) e-commerce platform in Fig. 1. The co-purchasing network may evolve due to the e-commerce platform being available in new regions. The evolution can lead to the introduction of new product categories and changes in the relative popularity of existing ones. It is crucial to identify products in potential novel categories due to either safety reasons or the need to develop new insights for platform improvement. Additionally, the detection has to be able to be done under the subpopulation shift among existing (non-novel) categories, ideally without knowing the category labels of individual products due to the chaotic nature of a c2c platform. Similar phenomena and demands also exist in academic citation graphs where papers demonstrating new research topics are to be detected. These applications underscore the importance of studying novel node category detection amidst subpopulation shifts.
@@ 2.1 PU Learning and Novel Category Detection
PU learning [1] is the problem of learning from a dataset with only positively labeled and unlabeled data. Novel category detection [3, 27] can be naturally reframed as a PU learning problem. The concept of novel and non-novel in our problem can be mapped to the concept of negative and positive in PU learning, respectively. Then, the role of the source nodes would correspond to the positively labeled samples in PU learning since all source nodes are known to be from the non-novel (positive) categories. The target nodes would correspond to the unlabeled samples because they could be from a novel (negative) or non-novel (positive) category and the learner does not know which ones are novel and which ones are not. This connection has been drawn by several prior work [3, 10, 27]. Mainstream PU learning methods [7, 8, 17, 38] deal with the absence of negative labels through risk estimator design, treating unlabeled samples as negatives and labeled samples as weighted combinations of positives and negatives. These risk estimators assume the class priors are given. In practice, the class priors have to be estimated from data by mixture proportion estimation (MPE) techniques [11, 35]. These PU learning approaches are based on the Selected Completely at Random (SCAR) assumption [8]. In our context, SCAR would assume that every node from a non-novel category has an equal probability of being in the source domain.
@@ 4.1 Recall-Constrained Optimization
Figure 2: An illustration of RECO-SLIP. The upper-right module is the recall-constrained optimization component (Eq. 3) where the classifier adjusts its scores to minimize the FPR on the source while reserving enough target nodes as novel. The bottom module is the selective link prediction component where the link prediction loss (Eq. 7) is imposed on the target subgraph excluding the nodes with the highest scores (Eq. 4, 5, 6). A solid orange arrow pair between two nodes denotes their representation similarity is maximized whereas a bidirectional hollow arrow denotes the similarity is minimized.
@@ 4.2 Selective Link Prediction
Using the domain labels as a proxy of novel/non-novel labels essentially views all target nodes as equally novel. The novel and non-novel node representations within the target domain will not be sufficiently separated since the domain labels confuse them as the same. However, the edge connection pattern can reveal additional information due to the homophily property. For instance, the target novel nodes should have few edge connections to the target non-novel nodes since they belong to different categories. Therefore, we use link prediction with the graph autoencoder (GAE) [15] objective as an auxiliary task during training to preserve the category subgroup structure in the node representation space.
@@
$\mathcal{L}_{lp}(g)=-\frac{1}{|\mathcal{E}_{+}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{+}}\log\sigma(\mathbf{g}_{v_{i}}\cdot\mathbf{g}_{v_{j}})-\frac{1}{|\mathcal{E}_{-}^{*}|}\sum_{(v_{i},v_{j})\in\mathcal{E}_{-}^{*}}\log(1-\sigma(\mathbf{g}_{v_{i}}\cdot\mathbf{g}_{v_{j}}))$ (7)
@@ 5.1.1 Data
We evaluate RECO-SLIP and baseline methods on five widely used public benchmark datasets: Cora [24], CiteSeer [24], Computers [25], Photo [25], and arxiv [13]. Cora, CiteSeer, and arxiv are academic citation graphs while Computers and Photo are product co-purchasing networks. We use the original labels provided by the datasets as the category labels. For Cora, CiteSeer, Computers, and Photo, we view the last class as the novel category and simulate subpopulation shifts by splitting nodes in each category into source and target based on Table 1. We name the preprocessed versions of the four datasets with “-S” at the end representing “shift”. arxiv is a subset of the ogbn-arxiv graph [13] containing $6$ robotics-related categories from $1990$ to $2012$. arxiv enables a natural source/target split and a novel category emergence since each node is associated with a timestamp and nodes belonging to the category “cs.SY: systems and control” did not exist until $2007$. Therefore, we define nodes with timestamps before $2007$ as the source and the others as the target. In terms of train/test data split, we consider the commonly used transductive node classification setup. The test nodes of arxiv are the nodes with timestamps in $2012$ while the test nodes of the other four datasets are randomly selected from the target domain. We further split a part of the training nodes into the validation set for model selection. We show the dataset-specific statistics in Table 2.
@@ 5.1.2 Baselines
  • SAR-EM [2]: SAR-EM is a propensity-weighting approach for PU-learning when SCAR does not hold. It jointly learns the PU classifier and the propensity score of each data point via the EM algorithm. We select this method as the representative method of propensity-weighting approaches due to its relatively flexible assumptions and publicly available code implementation. Later work using propensity-weighting [12, 28] imposes strong assumptions such as Local Certainty and Probabilistic Gap [^3] Local Certainty assumes the relationship between the observed features and the true class is a deterministic function, meaning that the class distributions do not overlap. Probabilistic Gap allows overlapping class distributions but assumes that the propensity scores follow the ordering of the posterior probabilities, i.e. $p(y=1|\mathbf{x})$. and do not provide code. Therefore, we omit these methods in our experiments.
@@ 5.1.3 Hyper-Parameters
All methods except LP-PUL use a two-layer graph convolution network (GCN) [16] coupled with a two-layer MLP as the classifier architecture. We find that the Adam optimizer [14] with a learning rate of $0.001$ works well with the domain discriminator components in all methods except LP-PUL. This optimizer setting also works well for the primal and dual optimizers of RECO-SLIP. For hyper-parameters specific to RECO-SLIP, we set $\xi=0.001$, $T=1000$, and $\tilde{\bm{\alpha}}=[0.05,0.1,0.15,0.2,0.25]$ for all datasets. By default, we set $\tilde{\beta}$ to $0.01$. However, since the empirical FPR estimate under all $\tilde{\alpha}$ could not go below $0.01$ on Photo-S and CiteSeer-S due to dataset characteristics, we set $\tilde{\beta}$ to $0.05$ for these two datasets.